OpenStudy (anonymous):

what common fraction is one third of the way from 1/4 to 1/3?

OpenStudy (anonymous):

The distance from 1/4 to 1/3 is \[\frac{1}{3} - \frac{1}{4} = \frac{4}{12} - \frac{3}{12} = \frac{1}{12} \] One third of this is \[\frac{1}{3} \frac{1}{12} = \frac{1}{36} \] So, if you add that to 1/4, \[\frac{1}{4} + \frac{1}{36} = \frac{9}{36} + \frac{1}{36} = \frac{10}{36} = \frac{5}{18} \]

OpenStudy (anonymous):

thanx